how would i solve this equation.|4-3x|=|2-3x| . i know the answer is 1. i just dont know how to get there
Answer by josgarithmetic(39838)

Without getting into too much detail, either 3x-4=2-3x or 4-3x=2-3x. Maybe both ways will work? Solve each and find if one or both will give you solutions, maybe different solutions.

If both expressions inside abs val function are positive or negative, then just remove the abs value markers. If either expression is positive but the other negative, then reverse ONE of the expressions and keep the other one the same, and then remove the absolute value markers.